APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Wednesday December 12, 2007 (MJD 54447) ---=== OBSERVING TEAM ===--- Elena Malanushenko (day) Stephanie Snedden (observer) Shannon Watters (observer) ---=== OBSERVING PLAN ===--- Primary mirror collimation, pointing model, science. ---=== OBSERVING SUMMARY ===--- The dew point depression was < 4C at the beginning of the night. Eventually it increased enough for us to start ventilating and open. When we tried to move the enclosure it was moving sporadically and the warning bell kept going on and off (it usually just 'rings' once at the beginning of movement). It moved 3-4 feet and then stopped. French was here and found that there was no power to the rails because ice had built up on them. We used the winches to get the enclosure back over the telescope and waited for the rail heater to melt the ice. After 2 hours of waiting we tried moving the enclosure again but it stopped at the same spot and wouldn't move back. Since the weather wasn't threatening we left it in that position waiting for further defrosting. We thought about waking French (08:15Z) to help move the enclosure so we could get on sky but the seeing on the 3.5m was deteriorating to 1.7" so we would have ~2.0" which wouldn't be useful for collimation. French and crew will work on the enclosure/rails this morning. We had to bypass the Bldg_Clear_Alt interlock to move the telescope to 30 alt while the enclosure is partly off so that bypass is still interlocked. Astrometric monitor #3 was flickering and smelled like burning plastic so we shut it off. We had a watcher warning for VS+ error on CCD55 but this is a warning we get when it's cold out and it went back into spec shortly. We had to run endNight manually starting with the copying of the IR camera images because it failed on the queryProcessMtData stage (we think because there was a missing MT file number). ---=== OBSERVING LOG ===--- Afternoon: --------- French and DanL worked with mirror. Limited check of equipment have been performed. I have to reboot photo crates as id4 did not respond. It was nessasary to clear the pool as it does not have more room availble (initNight -poolInit). The new az fiducial table has been checked in until to step of switching MCP. New version is v1_68. Nominal values of mirror number changed and the difference between real and nominal numbers are more then 0.05. Spectro collimation did not measured and set because we did not have spectro cart on telescope mounted. Now the equipment work properly. Imager is mounted. The humidity is 95% and no ventilation started. Night: ------ endNight finished (manually). ---=== IMAGING RUN SUMMARY ===--- Run Time Stripe Lambda Last Flavor Comments Start End Begin End Frame ------------------------------------------------------------------------- 54446 22:12Z 22:14Z 100 O 123.33 123.94 47847 ignore 54446 22:12Z 22:14Z 100 O 123.33 123.94 47847 ignore 54446 22:16Z 22:23Z 100 O 124.42 126.26 47847 ignore 7231 22:26Z 22:34Z 100 O 127.05 128.92 27 ignore 7232 22:41Z 23:15Z 100 O 130.64 139.31 68 ignore ---=== IMAGING RUN DETAILS ===--- ---=== SKIPPY RESULTS ===--- Run Frame nFrames stars muErr muRms nuErr nuRms rot az el --------------------------------------------------------------------------- ---=== LTMATCH RESULTS ===--- Run Field nFields alt az nGood rowMean rowSig colMean colSig rot ------------------------------------------------------------------------ ---=== SPECTROSCOPY DATA SUMMARY ===--- Summary Checked (y/n): QA Procedures Done (y/n): UT Exp Time flavor comment (S/N)^2 totals ========================================== b1 r1 b2 r2 ----- sequence 47848, plate -9999 ------- 22:54 47848 0.0 bias ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Instrument Az Alt Rot Scale pos offset pos offset pos offset ------------------------------------------------------------------------------ ---=== TELESCOPE OFFSETS AND SCALE II ===--- ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm ------------------------------------------------------------------------------ ---=== WEATHER LOG ===--- Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky 21:03Z 23 19 4 289 (WNW) 272 - 21:36Z 22 20 6 349 (NNW) 472 - 22:08Z 22 20 4 355 (NNW) 345 - 22:39Z 22 19 4 358 (NNW) 239 - 23:10Z 23 18 6 306 (NW) 181 - 23:40Z 22 19 2 304 (NW) 215 - 00:13Z 21 20 0 296 (WNW) 279 - clear 00:45Z 21 18 14 9 (N) 258 - 01:17Z 22 11 2 336 (NNW) 85 - 01:49Z 21 16 8 333 (NNW) 143 - 02:25Z 21 16 6 348 (NNW) 116 - 02:58Z 21 17 7 358 (NNW) 136 - 03:30Z 22 11 5 12 (NNE) 68 - 04:02Z 22 14 12 20 (NNE) 91 - 04:34Z 22 8 12 28 (NNE) 224 - 05:06Z 24 0 9 4 (N) 80 - 05:38Z 25 -2 8 357 (NNW) 96 - 06:10Z 23 3 2 48 (NE) 361 - 06:43Z 22 3 16 6 (N) 62 - 07:15Z 23 -2 2 0 (N) 95 - 07:47Z 22 -3 12 34 (NE) 107 - 08:23Z 25 -11 11 40 (NE) 120 - 08:59Z 27 -13 9 39 (NE) 43 - 09:30Z 28 -12 12 2 (N) 92 - 10:03Z 30 -15 6 342 (NNW) 81 - 10:36Z 26 -7 9 359 (NNW) 71 - 11:07Z 27 -9 7 66 (ENE) 91 - 11:38Z 33 -13 7 346 (NNW) 68 - 12:08Z 32 -15 5 353 (NNW) 77 - clear 12:40Z 33 -14 4 285 (WNW) 65 - 13:12Z 34 -13 3 269 (W) 73 - ---=== TELESCOPE STATUS ===--- 01:00Z - Doors and louvers opened, fans on. 02:50Z - Doors closed, fans and blowers off. 03:20Z - Doors opened, fans and blowers on. 12:20Z - Fans and blowers off, louvers closed (doors open, enclosure partly off) Telescope stowed at 30 degrees with the counterweights set to 705 and the spectro cryo system is not connected (day crew working on enclosure problem). The Bldg_Clear_Alt interlock is still bypassed. 12:10Z Spectro dewar: 44.2 lb, 0.06 psi 12:10Z Imager dewar: 86.0 lb, 19.24 psi ---=== SOFTWARE USED ===--- IOP/SOP: v4_38_0 Watcher: v2_36_0 MCP: v6_4_0 TPM: tpm_v3_5_0 AstroDa: v15_13_1 TCC: v2.7.2.1 Aug 04 sdssProcedures: v2_37 SoS: v5_1_8 plate-mapper: v4_3_1 ---=== MIRROR NUMBERS ===--- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A -5.2150 -5.2130 Axial B -0.2124 -0.2135 Axial C -0.0017 -0.0016 Trans D 11.6430 11.6400 Lateral E 9.9695 9.9568 Lateral F 11.5443 11.5316 GALILS Commanded: 4600. -2000. 2650. -9300. -9800. -9800. Actual: 4600. -2000. 2650. -9300. -9800. -9800. SETMIR VALUES PriDesOrient: 0.00 0.00 0.00 0.00 0.00 PriOrient: 0.00 0.00 0.00 0.00 0.00 SECONDARY: ---------- Focus: 0.00 Air Temp.: -5.7 Alt.: 90.004470 MIGS TONIGHT NOMINAL Axial A 0.8529 0.7973 Axial B 0.6626 0.6041 Axial C 0.5982 0.5425 Trans D 0.6100 0.5999 GALILS Commanded: 44038. 69839. -113619. 4100. 4900. Actual: 43723. 69395. -112888. 4082. 4893. SETMIR VALUES SecDesOrient: 1257.00 60.00 -9.43 -50.00 240.90 SecOrient: 1256.94 59.95 -9.47 -49.92 240.46 ---=== PROBLEMS IN DETAIL ===--- No spectro bias done. No spectro collimation.